OpenStudy (anonymous):

what is the x values in y = –16x2 + 250x + 0

OpenStudy (cwrw238):

do you mean the zeros of -16x^2 + 250x = 0?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

factor that quadratic: -2x(8x - 125) = 0 x(8x - 125) = 0 can you take it from here?